OpenStudy (anonymous):

Find the prime factorization of the following integer: 270

OpenStudy (anonymous):

\[2 * 3^3 * 5\]

OpenStudy (anonymous):

2, 5, 3,3,3

OpenStudy (llort):

beat me to it, 2*3^3*5=270

OpenStudy (anonymous):

\[2 \times 5 \times 3^3\]

OpenStudy (anonymous):

2 * 3^3 * 5